# Distribution-Free Calibration of Statistical Confidence Sets

### v1

@@ Abstract
Constructing valid confidence sets is a crucial task in statistical inference, yet traditional methods often face challenges when dealing with complex models or limited observed sample sizes. These challenges are frequently encountered in modern applications, such as Likelihood-Free Inference (LFI). In these settings, confidence sets may fail to maintain a confidence level close to the nominal value. In this paper, we introduce two novel methods, TRUST and TRUST++, for calibrating confidence sets to achieve distribution-free conditional coverage. These methods rely entirely on simulated data from the statistical model to perform calibration. Leveraging insights from conformal prediction techniques adapted to the statistical inference context, our methods ensure both finite-sample local coverage and asymptotic conditional coverage as the number of simulations increases, even if $n$ is small. They effectively handle nuisance parameters and provide computationally efficient uncertainty quantification for the estimated confidence sets. This allows users to assess whether additional simulations are necessary for robust inference. Through theoretical analysis and experiments on models with both tractable and intractable likelihoods, we demonstrate that our methods outperform existing approaches, particularly in small-sample regimes. This work bridges the gap between conformal prediction and statistical inference, offering practical tools for constructing valid confidence sets in complex models.
Keywords: confidence sets, distribution-free inference, likelihood-free inference, nuisance parameters, uncertainty quantification, conformal prediction, finite-sample coverage
@@ 1.1.2 Likelihood-Free Inference
There is also a substantial body of work on Bayesian LFI, but these methods do not achieve conditional coverage (Eq. 1) even as $B\to\infty$ (Hermans et al., 2021).
@@ 1.1.3 Conformal Prediction
The first step in a standard conformal method involves creating a conformity score $\hat{s}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which measures how well the outputs of a regression model predict new labels $Y\in\mathcal{Y}$. In a regression setting, a standard choice for the conformity score is the regression residual, given by $\hat{s}(\mathbf{x},y)=|y-\widehat{\mathbb{E}}[Y|\mathbf{x}]|$. The prediction region then takes the form $R(\mathbf{x})=\{y:\hat{s}(\mathbf{x},y)\leq t\}$, where $t$ is chosen to be the $(1-\alpha)$-quantile of the residuals $\hat{s}(\mathbf{x},y)$ evaluated on a holdout set not used to train $\hat{s}$.
@@
Another key difference between prediction intervals and confidence sets is that $\theta$ is not random, whereas $Y$ is. Thus, marginal coverage (Eq. 3) is insufficient for statistical inference problems; typically, coverage for all fixed $\theta$’s, corresponding to conditional coverage (Eq. 1), is also desired. This is because the distribution $r(\theta)$, used to sample $\theta$ on the training set, is often arbitrary; it does not encode any randomness observed in the real world. Unfortunately, it is well-known that conditional coverage can’t be achieved without strong assumptions (Lei and Wasserman, 2014). While many conformal approaches aim to achieve conditional coverage asymptotically, they only control coverage conditional on the features $\mathbf{x}$, and not $y$. The exception to this are label-conditional conformal prediction methods (Vovk et al., 2014, 2016; Sadinle et al., 2019), but these apply only to classification problems.
@@ 1.1.4 Partition-Based Conformal Methods
Our approach to asymptotically controlling conditional coverage is closely related to the works of Vovk et al. (2005); Lei and Wasserman (2014); Boström and Johansson (2020); Izbicki et al. (2020, 2022). These methods first partition the feature space and then calibrate $t$ using the conformal approach described in the previous section for each partition element separately. In particular, TRUST is closely related to LoCart (Cabezas et al., 2024), which creates a data-driven partition of $\mathcal{X}$ using regression trees designed to ensure that
@@ 1.1.5 Conformal Predictive Distributions
In simple terms, $Q$ provides an estimated probability distribution for the test labels $y$ based on conformity scores. Specifically, if the data is i.i.d., $Q$ is valid, meaning that $Q((\mathbf{X}_{n+1},y),u)$ follows a uniform distribution (Vovk et al., 2019), and thus $Q$ is a valid cumulative distribution function. Leveraging the validity of $Q$, we construct $1-\alpha$ confidence CP prediction intervals by utilizing the specific $\alpha/2$ and $21-\alpha/$ 2 percentiles of the CPD (Boström et al., 2021).
@@ 2 Methodology
$\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta).$
@@ 2.1.1 Partition-based Estimate
$\widehat{H}_{B}(t|\theta,\mathcal{A}):=\frac{1}{|I_{A(\theta)}|}\sum_{b\in I_{A(\theta)}}\mathbb{I}\left(\tau(\mathbf{X}_{b},\theta_{b})\leq t\right),$ (5)
@@
With this partition-based construction of $\widehat{H}$, the plugin cutoff $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$, used to estimate confidence sets, corresponds to the $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$. Also, $\widehat{H}_{B}$ is essentially a conditional predictive conformal distribution (Section 1.1.5), with the exception that we do not add the uniform distribution used to ensure continuity as it is not needed here. This method therefore aligns with a partition-based conformal approach. However, unlike standard partition-based conformal methods, the partition $\mathcal{A}$ must be selected to approximate $H(\alpha|\theta)$ well. The next section discusses how to choose $\mathcal{A}$ to achieve this goal.
@@ 2.1.2 Choosing the partition: TRUST
       $\widehat{C}_{\theta,B}\leftarrow\alpha$-quantile of $\{\tau(X_{c},\theta_{c})\}_{c\in J_{A(\theta)}}$;
@@
Figure 2: Schematic flowchart of TRUST. Over a simulated dataset $(\mathbf{X}_{1},\theta_{1}),\dots,(\mathbf{X}_{B},\theta_{B})$, we compute the statistic $\tau$ and create the set $\mathcal{I}$, which is splitted between $\mathcal{I}_{train}$ and $\mathcal{I}_{cal}$. $\mathcal{I}_{train}$ is used to train a regression tree that predicts $\tau(\mathbf{X},\theta)$ using $\theta$. The tree partitions the parametric space in $A_{1},\dots,A_{k}$. If a $\theta^{*}$, from the grid, falls inside the $A_{1}$ partition then, in order to calculate the cutoff $\widehat{C}_{\theta^{*},B}$, we will compute the $\alpha$-quantile of all $\tau(\mathbf{X},\theta)$’s from which $\theta$ belongs both to $\mathcal{I}_{cal}$ and $A_{1}$. Finally, we use the calculated cutoff to find the confidence set $\widehat{R}_{B}$.
@@ 2.2 From trees to forests: TRUST++
Since a single regression tree often lacks expressiveness, we extend our approach to random forests. However, averaging the $\widehat{C}_{\theta,B}$ values obtained across trees does not ensure distribution-free guarantees (Cabezas et al., 2024). Instead, we use a random forest to design a new partition of $\Theta$, and then apply the methodology described in Section 2.1.1. Specifically, we first create $K$ TRUST regression trees, each on a different bootstrap sample of the simulated dataset $\{(\theta_{1},\mathbf{X}_{1}),\ldots,(\theta_{B},\mathbf{X}_{B})\}$. Next, let $\rho(\theta^{\prime},\theta)$ denote Breiman’s proximity measure (Breiman, 2001), which counts the number of times $\theta^{\prime}$ and $\theta$ appear in the same leaf across the $T$ trees. Finally, we define the partition $\mathcal{A}$ as the partition induced by the equivalence relation
@@
       $\widehat{C}_{\theta,B}\leftarrow\alpha$-quantile of $\{\tau(\mathbf{X}_{c},\theta_{c})\}_{c\in J_{A(\theta)}}$;
@@ 2.4 Uncertainty About the Confidence Sets
Recall that, for each $\theta$, $C_{\theta}$ is estimated by computing the $\alpha$-quantile of the set $T_{A(\theta)}=\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$, while $C_{\theta}$ itself is the $\alpha$-quantile of the distribution of $\tau(\mathbf{X},\theta)|\theta$. Let $\tau_{A(\theta)}^{(i)}$ be the $i$-th order statistics of $T_{A(\theta)}$ and define $Z:=|\{\delta\in T_{A(\theta)}:\delta\leq C_{\theta}\}|$, where $Z$ represents the number of statistics in $T_{A(\theta)}$ that are less than or equal to the cutoff $C_{\theta}$. Note that if the statistics in $T_{A(\theta)}$ are identically distributed, then $Z\sim\text{Binomial}(n,1-\beta)$. Using this result, we adapt the method by Hahn and Meeker (2011) for constructing confidence intervals for quantiles to create a $1-\beta$ confidence set for $C_{\theta}$. This confidence set is defined as:
@@
where $\mathcal{I}(\mathbf{X})$ represents the set of parameter values confidently inside the confidence interval, $\mathcal{O}(\mathbf{X})$ contains values confidently outside, and $\mathcal{U}(\mathbf{X})$ includes those where the status remains uncertain. In the terminology of 3-way hypothesis testing (Berg, 2004; Esteves et al., 2016; Izbicki et al., 2023), $\mathcal{I}(\mathbf{X})$ is the acceptance region, $\mathcal{O}(\mathbf{X})$ is the rejection region, and $\mathcal{U}(\mathbf{X})$ is the agnostic region. The size of $\mathcal{U}(\mathbf{X})$ can be decreased by increasing the number of simulations used to estimate $C_{\theta}$, $B$. The next theorem shows that this approach controls the probability of incorrect conclusions.
@@
This method can be readily extended to the setting with nuisance parameters. Let $\eta^{*}$ be the configuration of nuisance parameters for which $\widehat{C}_{\mu,B}$ is computed (Equation 7). By construction, $\widehat{C}_{\mu,B}$ is the $\alpha$-quantile of the set $T_{A(\mu,\eta^{*})}=\{\tau(\mathbf{X}_{b},\mu_{b},\eta_{b}):b\in I_{A(\mu,\eta^{*})}\}$. Thus, we can derive a confidence set for the optimal cutoff using the same approach as in Eq. 8, based now on the set $T_{A(\mu,\eta^{*})}$.
@@ 3.1 Partition Coverage Guarantees
where $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$ represents the $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$, conditioned on $\theta$ belonging to the partition element $A(\theta)$. Then, we have
@@
which suggests that, with these methods, the probability that $\theta$ lies within the estimated region $\widehat{R}_{B}(\mathbf{X})$, given its association with $A(\theta)$, closely aligns with the nominal confidence level $(1-\alpha)$. In the following section, we will state the theorem formally and provide additional details surrounding this result.
Beyond the conditional coverage established above, these results also imply marginal coverage, ensuring that the confidence set $\widehat{R}_{B}(\mathbf{X})$ maintains the desired overall coverage probability:
@@
$\mathbb{P}(\theta\in\widehat{R}_{B}(\mathbf{X}))\geq 1-\alpha,$
ensuring the marginal coverage of the constructed confidence set. In particular, the partition can be taken as that provided by TRUST or TRUST++.
@@ 3.2 TRUST and TRUST++ Conditional Coverage Guarantees
where $\widehat{C}_{\theta,B}$ is the cutoff determined by the partition, representing the $\alpha$-quantile of the distribution of $\tau(\mathbf{X}_{b},\theta_{b})$ values within the partition containing $\theta$.
Then, if Assumption 1 holds, both TRUST and TRUST++ are asymptotically consistent, that is:
@@ 4 Implementation Details and Tuning Parameters
To address this, we propose a straightforward grid-search algorithm for optimizing $M$ using a small, additional simulated validation grid. The main idea is to select $M$ from a fine grid between 00 and $K$ that minimizes estimated deviation from conditional coverage. This is done by calculating coverage across the validation grid with a batch of statistics simulated at each fixed grid point and then computing the mean absolute error of the estimated coverage relative to the nominal confidence level $1-\alpha$. Section 5 provides further details on this process. Alternatively, coverage can be estimated using the LF2I diagnostic module (Dalmasso et al., 2021), which also leverages an additional simulation set. Algorithm 3 in Appendix B outlines the tuning procedure for $M$.

### v2

# Conformal Calibration of Statistical Confidence Sets

@@ Abstract
Constructing valid confidence sets is a crucial task in statistical inference, yet traditional methods often face challenges when dealing with complex models or limited observed sample sizes. These challenges are frequently encountered in modern applications, such as Likelihood-Free Inference (LFI). In these settings, confidence sets may fail to maintain a confidence level close to the nominal value. In this paper, we introduce two novel methods, TRUST and TRUST++, for calibrating confidence sets to achieve distribution-free conditional coverage. These methods rely entirely on simulated data from the statistical model to perform calibration. Leveraging insights from conformal prediction techniques adapted to the statistical inference context, our methods ensure both finite-sample local coverage and asymptotic conditional coverage as the number of simulations increases, even if $n$ is small. They effectively handle nuisance parameters and provide computationally efficient uncertainty quantification for the estimated confidence sets. This allows users to assess whether additional simulations are necessary for robust inference. Through theoretical analysis and experiments on models with tractable and intractable likelihoods, we demonstrate that our methods outperform existing approaches, particularly in small-sample regimes. This work bridges the gap between conformal prediction and statistical inference, offering practical tools for constructing valid confidence sets in complex models.
@@ 1.1.2 Likelihood-Free Inference
There is also a substantial body of work on Bayesian LFI, but these methods do not achieve conditional in the sense of Eq. 1) even as $B\to\infty$ (Hermans et al., 2021).
@@ 1.1.3 Conformal Prediction
The first step in a standard conformal method involves creating a conformity score $\hat{s}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, which measures how well the outputs of a regression model predict new labels $Y\in\mathcal{Y}$. In a regression setting, a standard choice for the conformity score is the regression residual, given by $\hat{s}(\mathbf{x},y)=|y-\widehat{\mathbb{E}}[Y|\mathbf{x}]|$. The prediction region then takes the form $R(\mathbf{x})=\{y:\hat{s}(\mathbf{x},y)\leq t\}$, where $t$ is chosen to be the adjusted $(1-\alpha)$-quantile of the residuals $\hat{s}(\mathbf{x},y)$ evaluated on a holdout set not used to train $\hat{s}$.
@@
Another key difference between prediction intervals and confidence sets is that $\theta$ is not random, whereas $Y$ is. Thus, marginal coverage (Eq. 3) is insufficient for statistical inference problems; typically, coverage for all fixed $\theta$’s, corresponding to conditional coverage (Eq. 1), is also desired. This is because the distribution $r(\theta)$, used to sample $\theta$ on the training set, is often arbitrary; it does not encode any randomness observed in the real world. Unfortunately, it is well-known that conditional coverage can’t be achieved without strong assumptions (Lei and Wasserman, 2014). While many conformal approaches aim to achieve conditional coverage asymptotically, they only control coverage conditional on the features $\mathbf{x}$, and not $y$. The exception to this are label-conditional conformal prediction methods (Vovk et al., 2014, 2016; Sadinle et al., 2019; Ding et al., 2023), but these apply only to classification problems.
@@ 1.1.4 Partition-Based Conformal Methods
Our approach to asymptotically controlling conditional coverage is closely related to the works of Vovk et al. (2005); Lei and Wasserman (2014); Boström and Johansson (2020); Izbicki et al. (2020, 2022); Vovk et al. (2022). These methods first partition the feature space and then calibrate $t$ using the conformal approach described in the previous section for each partition element separately. In particular, TRUST is closely related to LoCart (Cabezas et al., 2025), which creates a data-driven partition of $\mathcal{X}$ using regression trees designed to ensure that
@@ 1.1.5 Conformal Predictive Distributions
In simple terms, $Q$ provides an estimated probability distribution for the test labels $y$ based on conformity scores. Specifically, if the data is i.i.d., $Q$ is valid, meaning that $Q((\mathbf{X}_{n+1},y),u)$ follows a uniform distribution (Vovk et al., 2019), and thus $Q$ is a valid cumulative distribution function. Leveraging the validity of $Q$, we construct $1-\alpha$ confidence CP prediction intervals by utilizing the specific $\alpha/2$ and $1-\alpha/2$ percentiles of the CPD (Boström et al., 2021).
@@ 2 Methodology
$\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta),$
where $\widehat{H}_{B}$ is the adjusted empirical distribution function defined as
$\widehat{H}_{B}(t|\theta):=\frac{1}{B+1}\left(\sum_{b=1}^{B}\mathbf{1}\{\tau(\mathbf{X}^{(b)},\theta)\leq t\}+1\right).$
@@ 2.1.1 Partition-based Estimate
$\widehat{H}_{B}(t|\theta,\mathcal{A}):=\frac{1}{|I_{A(\theta)}|+1}\left(\sum_{b\in I_{A(\theta)}}\mathbb{I}\left(\tau(\mathbf{X}_{b},\theta_{b})\leq t\right)+1\right),$ (5)
@@
With this partition-based construction of $\widehat{H}$, the plugin cutoff $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$, used to estimate confidence sets, corresponds to the adjusted $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$. Also, $\widehat{H}_{B}$ is essentially a conditional predictive conformal distribution (Section 1.1.5), with the exception that we do not add the uniform distribution used to ensure continuity as it is not needed here. This method therefore aligns with a partition-based conformal approach. However, unlike standard partition-based conformal methods, the partition $\mathcal{A}$ must be selected to approximate $H(\alpha|\theta)$ well. The next section discusses how to choose $\mathcal{A}$ to achieve this goal.
@@ 2.1.2 Choosing the partition: TRUST
 $\widehat{C}_{\theta,B}\leftarrow$ adjusted $\alpha$-quantile of $\{\tau(X_{c},\theta_{c})\}_{c\in J_{A(\theta)}}$;
@@
Figure 2: Schematic flowchart of TRUST. Over a simulated dataset $(\mathbf{X}_{1},\theta_{1}),\dots,(\mathbf{X}_{B},\theta_{B})$, we compute the statistic $\tau$ and create the set $\mathcal{I}$, which is splitted between $\mathcal{I}_{train}$ and $\mathcal{I}_{cal}$. $\mathcal{I}_{train}$ is used to train a regression tree that predicts $\tau(\mathbf{X},\theta)$ using $\theta$. The tree partitions the parametric space in $A_{1},\dots,A_{k}$. If a $\theta^{*}$, from the grid, falls inside the $A_{1}$ partition then, in order to calculate the cutoff $\widehat{C}_{\theta^{*},B}$, we will compute the adjusted $\alpha$-quantile of all $\tau(\mathbf{X},\theta)$’s from which $\theta$ belongs both to $\mathcal{I}_{cal}$ and $A_{1}$. Finally, we use the calculated cutoff to find the confidence set $\widehat{R}_{B}$.
@@ 2.2 From trees to forests: TRUST++
Since a single regression tree often lacks expressiveness, we extend our approach to random forests. However, averaging the $\widehat{C}_{\theta,B}$ values obtained across trees does not ensure distribution-free guarantees (Cabezas et al., 2025). Instead, we use a random forest to design a new partition of $\Theta$, and then apply the methodology described in Section 2.1.1. Specifically, we first create $K$ TRUST regression trees, each on a different bootstrap sample of the simulated dataset $\{(\theta_{1},\mathbf{X}_{1}),\ldots,(\theta_{B},\mathbf{X}_{B})\}$. Next, let $\rho(\theta^{\prime},\theta)$ denote Breiman’s proximity measure (Breiman, 2001), which counts the number of times $\theta^{\prime}$ and $\theta$ appear in the same leaf across the $T$ trees. Finally, we define the partition $\mathcal{A}$ as the partition induced by the equivalence relation
@@
 $\widehat{C}_{\theta,B}\leftarrow$ adjusted $\alpha$-quantile of $\{\tau(\mathbf{X}_{c},\theta_{c})\}_{c\in J_{A(\theta)}}$;
@@ 2.4 Uncertainty About the Confidence Sets
Recall that, for each $\theta$, $C_{\theta}$ is estimated by computing the adjusted $\alpha$-quantile of the set $T_{A(\theta)}=\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$, while $C_{\theta}$ itself is the $\alpha$-quantile of the distribution of $\tau(\mathbf{X},\theta)|\theta$. Let $\tau_{A(\theta)}^{(i)}$ be the $i$-th order statistics of $T_{A(\theta)}$ and define $Z:=|\{\delta\in T_{A(\theta)}:\delta\leq C_{\theta}\}|$, where $Z$ represents the number of statistics in $T_{A(\theta)}$ that are less than or equal to the cutoff $C_{\theta}$. Note that if the statistics in $T_{A(\theta)}$ are identically distributed, then $Z\sim\text{Binomial}(n,1-\beta)$. Using this result, we adapt the method by Hahn and Meeker (2011) for constructing confidence intervals for quantiles to create a $1-\beta$ confidence set for $C_{\theta}$. This confidence set is defined as:
@@
where $\mathcal{I}(\mathbf{X})$ represents the set of parameter values confidently inside the confidence interval, $\mathcal{O}(\mathbf{X})$ contains values confidently outside, and $\mathcal{U}(\mathbf{X})$ includes those where the status remains uncertain. In the terminology of 3-way hypothesis testing (Berg, 2004; Esteves et al., 2016; Izbicki et al., 2025), $\mathcal{I}(\mathbf{X})$ is the acceptance region, $\mathcal{O}(\mathbf{X})$ is the rejection region, and $\mathcal{U}(\mathbf{X})$ is the agnostic region. The size of $\mathcal{U}(\mathbf{X})$ can be decreased by increasing the number of simulations used to estimate $C_{\theta}$, $B$. The next theorem shows that this approach controls the probability of incorrect conclusions.
@@
This method can be readily extended to the setting with nuisance parameters. Let $\eta^{*}$ be the configuration of nuisance parameters for which $\widehat{C}_{\mu,B}$ is computed (Equation 7). By construction, $\widehat{C}_{\mu,B}$ is the adjusted $\alpha$-quantile of the set $T_{A(\mu,\eta^{*})}=\{\tau(\mathbf{X}_{b},\mu_{b},\eta_{b}):b\in I_{A(\mu,\eta^{*})}\}$. Thus, we can derive a confidence set for the optimal cutoff using the same approach as in Eq. 8, based now on the set $T_{A(\mu,\eta^{*})}$.
@@ 3.1 Partition Coverage Guarantees
where $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$ represents the adjusted $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$, conditioned on $\theta$ belonging to the partition element $A(\theta)$. Then, we have
@@
which suggests that, with these methods, the probability that $\theta$ lies within the estimated region $\widehat{R}_{B}(\mathbf{X})$, given its association with $A(\theta)$, closely aligns with the nominal confidence level $(1-\alpha)$. In the following section, we will state the theorem formally. Also, notice that although the local guarantee depends on the choice of $r(\theta)$ (and thus has a Bayesian flavor), the conditional one does not.
Beyond the conditional coverage established above, these results also imply marginal coverage:
@@
Let $\{(\theta_{1},\mathbf{X}_{1}),\dots,(\theta_{B},\mathbf{X}_{B})\}$ be an i.i.d. simulated dataset with a fixed partition $\mathcal{A}$ of $\Theta$, where each $\theta_{b}$ is drawn from a reference distribution $r(\theta)$ and each $\mathbf{X}_{b}$ is generated according to the statistical model with parameters $\theta_{b}$. For a test statistic $\tau(\mathbf{X},\theta)$, consider the confidence set constructed by our methods. Then $\mathbb{P}(\theta\in\widehat{R}_{B}(\mathbf{X}))\geq 1-\alpha.$
@@ 3.2 TRUST and TRUST++ Conditional Coverage Guarantees
where $\widehat{C}_{\theta,B}$ is the cutoff determined by the partition, representing the adjusted $\alpha$-quantile of the distribution of $\tau(\mathbf{X}_{b},\theta_{b})$ values within the partition containing $\theta$. Then, if Assumption 1 holds, both TRUST and TRUST++ are asymptotically consistent, that is:
@@ 4 Implementation Details and Tuning Parameters
To address this, we propose a straightforward grid-search algorithm for optimizing $M$ using a small, additional simulated validation grid. The main idea is to select $M$ from a fine grid between 0 and $K$ that minimizes estimated deviation from conditional coverage. This is done by calculating coverage across the validation grid with a batch of statistics simulated at each fixed grid point and then computing the mean absolute error of the estimated coverage relative to the nominal confidence level $1-\alpha$. Section 5 provides further details on this process. Alternatively, coverage can be estimated using the LF2I diagnostic module (Dalmasso et al., 2021), which also leverages an additional simulation set. Algorithm 3 in Appendix B outlines the tuning procedure for $M$.